how many ounces of a 15% alcohol solution must be mixed with 9 ounces of a 20% alcohol solution to make a 18% alcohol solution
Answer by mananth(16946)

--- percent ---------------- Amount
Alcohol Solution I 20 ---------------- 9 ounces
Alcohol solution II 15 ---------------- x ounces
Mixture 18 ---------------- 9 + x ounces

20 * 9 + 15 x = 18 ( 9 + x )
180 + 15 x = 162 + 18 x
15 x -18 x = -180 + 162
-3 x = -18
/ -3
x = 6 ounces Alcohol solution II

m.ananth@hotmail.ca